# tableReferences for Table S1

### v1

@@ 4.2 Chest X-ray image generation
Figure 11: Comparing generated and real X-ray images. (a) Sketch of the embedded distributions of X-ray images from the three different datasets: real dataset (black), Progressive Growing Generative Adversarial Network (PGGAN) (grey), and Stable Diffusion (SD) model (gold). (b) Examples from real and generated X-ray images. Three full-view examples from each distribution and four examples magnifying the top right corner. (c) SW distance, C2ST, MMD (bandwidth=50), and FID between samples of the three distributions of embedded X-ray images. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
@@
In the second application we turned to a high dimensional example, in which we compared synthetic X-ray images generated by a Progressive Growing Generative Adversarial Network (PGGAN) model  (Segal et al., 2021) and by a StableDiffusion (SD) model (Malik & Humair, 2023) to real chest X-ray images from the ChestX-ray14 dataset (Wang et al., 2017). Each image has a total dimension of $1024\times 1024$ pixels.
@@
From visual inspection, we note two observations: First, the images produced by the SD model are clearer and sharper than either the real images or those generated by the PGGAN. Second, generated images contain unrealistic artifacts that distinguish them from real X-ray images (Fig. 11b). For instance, in real images, the top often contains annotations including e.g. patient id, side of the body, or the date the X-ray was taken. These textual elements often contain artifacts or, in case of SD, are completely unrealistic. To compare these high-dimensional images, we embedded them in a 512-dimensional embedding space using the CheXzero network (Tiu et al., 2022), a CLIP (Radford et al., 2021) network fine-tuned for chest X-ray images. We opted for using this specialized network instead of the standard InceptionV3 network as it might overcome biases introduced by classification task training (Kynkäänniemi et al., 2022). As expected, samples generated by PGGAN are closer to the real data across all distances compared to SD-generated data (Fig. 11c), likely due to the unrealistic sharpness and more obvious textual artifacts of the SD-generated images. However, C2ST is even high between PGGAN outputs and the real data, suggesting that the high-dimensionality of the data increases the sensitivity of this measure. Taken together, our results suggest that PGGAN is more accurate in generating realistic X-ray images compared to SD.
@@
Our findings highlight that using different metrics can support different conclusions. For instance, C2ST suggests equality between DDM1 and real decision time data, whereas SW distance and MMD metrics indicate a larger difference between DDM1 and the real data. Similarly, in analyzing X-ray image generation, SW distance, MMD, and FID metrics suggest a high similarity between PGGAN-generated and real images, whereas C2ST indicates a strong difference. Thus, we want to highlight the importance of using multiple complementary distances for best results and understanding of model limitations.
@@ 5 Discussion
This work describes and explores four commonly applied sample-based distances representing different methodologies for defining statistical distance: Using low-dimensional projections (SW), obtaining a distance using classifiers (C2ST), using embeddings through kernels (MMD) or neural networks (FID). Despite their operational differences, they are all based on a fundamental concept: simplifying complex distributions into more manageable feature representations to facilitate comparison. Sliced distances effectively reduce multidimensional distributions to a set of one-dimensional distributions, where classical metrics are more easily applied or calculated. MMD uses kernels to (implicitly) project samples into a higher dimensional feature space, in which comparing mean values becomes more expressive. Classifier-based methods (C2ST) transform the task of distribution comparison into a classification problem; comparison is made by investigating how well a classifier can distinguish the distributions. Lastly, network-based distances, such as FID, explicitly map samples into a representative feature space and compare distributions directly within this space.
@@
Sliced distances stand out for their computational efficiency in evaluating distributional discrepancies. However, when distributions differ primarily in lower-dimensional subspaces, sliced distances might not detect these subtle differences without a large number of slices (see Fig. 8c). There are approaches to reduce this effect by considering other projections than simple linear slices, as described in Section 2.1. Due to its computational efficiency and differentiability, the SW distance can also be used as a loss function to train generative models (Wu et al., 2019, Deshpande et al., 2018; 2019, Liutkus et al., 2019, Vetter et al., 2024). In our experiments, the metric did show convincing results and in contrast to the MMD, C2ST, and FID, SW distance does not require to choose specific hyperparameters for which results can differ drastically. Although currently not extensively used in literature for evaluation, this makes the SW distance efficient, scalable, and an objective baseline for general distribution comparisons. Yet, this also makes it less flexible to adapt to specific features of interest. Although, the majority of research on sliced distances focus on sliced Wasserstein metrics, slicing other metrics is also possible. For a certain subset of choices, equivalence to MMDs can be established (Kolouri et al., 2019).
@@
The Maximum Mean Discrepancy is a strong tool for comparing two groups of data by looking at their average values in a special feature space. The effectiveness of MMD largely depends on the kernel function chosen (implicitly representing the feature space), which affects how well it can spot differences between various types of data. Inappropriate kernel choice can leave the metric insensitive to subtle differences in the distribution (Gretton et al., 2012b, Sriperumbudur et al., 2009) (see Fig. 8). The MMD can be estimated efficiently and is differentiable, and thus often used as a loss function for training generative models (Arbel et al., 2019, Li et al., 2017, Bińkowski et al., 2021, Briol et al., 2019). Yet, a kernel must satisfy certain criteria, e.g., positive definiteness, making the design of new kernel functions challenging. Such constraints are relaxed for FID-like metrics, which focus on explicit representations of the embedding, whereas (kernel) MMD instead focuses on implicit representations. One advantage, however, is that the implicit embedding allows for infinite dimensional feature spaces (through characteristic kernel functions). These can be proven to be able to discriminate any two distinct distributions, something that is impossible through explicit representations used by the FID. Recently,  Kübler et al. (2022a) proposed a method to estimate MMD via a witness function that determines MMD (Appendix A.2). This method is closely related to C2ST in that both two estimate a discrepancy among distributions via a classifier (Kübler et al., 2022a, Section 5).
@@
Network-based approaches for evaluating distributions focus on the analysis of complex data, emphasizing the importance of capturing high-level, semantically meaningful features. These methods leverage neural networks to project data into a lower-dimensional, feature-rich space where traditional statistical distances can be applied more effectively. This is particularly important for tasks where the visual or semantic quality of the data is important, making them a popular choice for assessing generative models in domains such as image and text generation. The primary challenge lies in the design of suitable network architectures that can extract relevant features for accurate distribution comparison. Even more important than for the C2ST, this network must be well-established and shared which is a necessary but not sufficient criterion (Chong & Forsyth, 2020) to compare different results. While such well-established defaults exist for images (Szegedy et al., 2015a, Radford et al., 2021), this is not the case for other domains. For example, the time series generation community did not yet establish a default, and embedding nets are either trained or chosen by the authors (Smith & Smith, 2020, Jeha et al., 2021). We showed that the class-sensitivity of FID (Section 3.3) tends to model collapse (such as GANs), but might not necessarily reflect general image quality. In fact, Betzalel et al. (2022), Kynkäänniemi et al. (2022), Jayasumana et al. (2023) found that relevant features sometimes can disagree with human judgment and that CLIP embeddings align more closely to what humans perceive as favorable or unfavorable. Yet, FID features have been shown to align much better with human perception than traditional metrics (Zhang et al., 2018).
@@
Ultimately, the choice of distance hinges on the nature of the data under consideration and the specific characteristics of it one aims to compare. Given a specific dataset and problem, one will likely have to look beyond the distances discussed in this paper. For example, in the realm of human-centric data like images and audio, the perceptual indistinguishability of distributions is important (Gerhard et al., 2013, Zhang et al., 2018). Time series data, characterized by its temporal structure, demand metrics that account for temporal shifts and variations in a manner that does not disproportionately penalize minor discrepancies in timing, such as Dynamic Time Warping (Müller, 2007), or by using frequency information (Hess et al., 2023). In general however, irrespective of the specific use-case, it is advisable to use multiple distances in order to obtain a full picture, as using a single distances individually could support competing conclusions about the model that is to be evaluated.
Throughout this paper we have explained and analyzed four approaches of measuring statistical distance. While this is only a small subset of all possible distances available, we hope to have provided the foundational knowledge with which researchers can find, understand and interpret statistical distances specific to their own scientific application.
@@
Writing — Review & Editing: AD, AS, CS, FP, GM, JHM, JKL, JV, LH, LU, MG, MP, SB, ST, ZS
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039; DeepHumanVision, FKZ: 031L0197B; Simalesam: FKZ 01|S21055A-B), the Carl Zeiss Foundation (Certification and Foundations of Safe Machine Learning Systems in Healthcare), the Else Kröner Fresenius Stiftung (Project ClinbrAIn), and the European Union (ERC, DeepCoMechTome, 101089288). ST was partially supported by Grant-in-Aid for Research Activity Start-up 23K19966 and JST CREST JPMJCR2015. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SB, MD, MG, JK, JL, GM, MP, RR, AS, ZS and JV.
@@ A.1 Generative Models in Science
Table S1: Example generative models in science. Simulators include mechanistic models, DNN based models e.g. VAEs, GANs, Diffusion models. This is not an exhaustive list regarding disciplines using generative models nor generative models used in the listed disciplines.
ML model Simulator
natural science biology
@@ A.2 Details about Maximum Mean Discrepancy
Generally $\mathcal{X}$ and $\mathcal{H}$ are defined as a topological space, and the reproducing kernel Hilbert space (RKHS), respectively, but readers can simply think of the euclidean space $\mathbb{R}^{N}$ for the first examples in the main text.
@@
The definition of MMD can be rewritten through the notion kernel mean embedding. For given distribution $p(x)$, the kernel mean embedding $\mathbb{E}_{p(x)}[k(x,u)]\in\mathcal{H}$ is an element in RKHS that satisfies $\langle\mathbb{E}_{p(x)}[k(x,u)],f(u)\rangle_{\mathcal{H}}=\mathbb{E}_{p(x)}[f(x)]$ for any $f\in\mathcal{H}$ with argument $u\in\mathcal{X}$. The embedding $\mathbb{E}_{p(x)}[k(x,u)]$ is known to be determined uniquely if a corresponding kernel is bounded, i.e. $\|k(x,x^{\prime})\|_{\mathcal{H}}<\infty$ for any $x$. Then, as shown in Gretton et al. (2012a), $\mathsf{MMD}^{2}$ can be represented as
@@
A.3 Details about Wasserstein and Sliced-Wasserstein distances
Formally, the Wasserstein distance is described in measure-theoretic terms. We first state this definition, and then provide an accessible interpretation in the common case that the measures have well-defined probability density functions.
@@ A.3 Details about Wasserstein and Sliced-Wasserstein distances
Consider the case where $M\subseteq\mathbb{R}^{d}$, $\rho$ is the $q$-norm $||\cdot||_{q}$, and $\mu$ and $\nu$ have well-defined probability density functions $p_{1}$ and $p_{2}$ respectively. Then the Wasserstein-$q$ norm can be written as
$W_{q}(p_{1},p_{2})=\inf_{\gamma\sim\Gamma(p_{1},p_{2})}\left(\mathbb{E}_{x_{1},x_{2}\sim\gamma}\,||x_{1}-x_{2}||_{q}^{q}\right)^{\frac{1}{q}},$
where $\Gamma(p_{1},p_{2})$ is the set of all couplings, that is all possible “transportation plans”, between $p_{1}$, $p_{2}$. $\gamma\in\Gamma(p_{1},p_{2})$ is a joint distribution over $(x_{1},x_{2})$ with respective marginals $p_{1}$ and $p_{2}$ over $x_{1}$ and $x_{2}$.
The Sliced-Wasserstein distance is similarly defined in measure-theoretic terms for the measures $\mu$ and $\nu$. We refer the reader to Nadjahi (2021) for details. In the less general case described above, we can similarly provide a more intuitive definition. In particular, the random projection directions described in 2.1 are uniformly random vectors $u\in\mathbb{S}^{d-1}$, the unit sphere in $\mathbb{R}^{d}$. Projecting the distributions $p_{1}$ and $p_{2}$ onto $u$ induces one-dimensional distributions $p_{1}^{u}$ and $p_{2}^{u}$ with samples $u^{\mathsf{T}}x_{i}$, where $x_{1}\sim p_{1}$ and $x_{2}\sim p_{2}$. The Sliced-Wasserstein-$q$ distance can then be written as
$SW_{q}(p_{1},p_{2})=\mathbb{E}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{q}(p_{1}^{u},p_{2}^{u})],$
where $\mathcal{U}(\mathbb{S}^{d-1})$ is the uniform distribution over vectors on the unit sphere $\mathbb{S}^{d-1}$.
A.4 C2ST scores below 0.5

### v2

@@ 4.2 Chest X-ray image generation
Figure 11: Comparing generated and real X-ray images. (a) Sketch of the embedded distributions of X-ray images from the three different datasets: test set of real dataset (black), Progressive Growing Generative Adversarial Network (PGGAN) (grey), and Stable Diffusion (SD) model (gold). (b) Examples from real and generated X-ray images. Three full-view examples from each distribution and four examples magnifying the top right corner. (c) SW distance, C2ST, MMD (bandwidth=50), and FID between samples of the real and generated distributions of embedded X-ray images. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
@@
In the second application we turned to a high-dimensional example, in which we compared synthetic X-ray images generated by a Progressive Growing Generative Adversarial Network (PGGAN) model (Segal et al., 2021) and by a StableDiffusion (SD) model (Malik & Humair, 2023) to real chest X-ray images from the ChestX-ray14 dataset (Wang et al., 2017). Each image has a total dimension of $1024\times 1024$ pixels.
@@
From visual inspection, we note two observations: First, the images produced by the SD model are clearer and sharper than either the real images or those generated by the PGGAN. Second, generated images contain unrealistic artifacts that distinguish them from real X-ray images (Fig. 11b). In real images, the top often contains annotations including, e.g., patient id, side of the body, or the date the X-ray was taken. These textual elements often contain artifacts or, in case of SD, are completely unrealistic. To compare these high-dimensional images, we embedded them in a 512-dimensional embedding space using the CheXzero network (Tiu et al., 2022), a CLIP (Radford et al., 2021) network fine-tuned for chest X-ray images. We opted for using this specialized network instead of the standard InceptionV3 network as it might overcome biases introduced by classification task training (Kynkäänniemi et al., 2023).
As expected, samples generated by PGGAN are closer to the real data across all distances compared to SD-generated data (Fig. 11c), likely due to unrealistic sharpness and more obvious textual artifacts of the SD-generated images.However, C2ST is even high between PGGAN outputs and the real data, suggesting that the high-dimensionality of the data increases the sensitivity of this measure. Taken together, our results suggest that PGGAN is more accurate in generating realistic X-ray images compared to SD.
@@
Our findings show that using different metrics can support different conclusions. For instance, C2ST suggests equality between DDM1 and real decision time data, whereas SW distance and MMD metrics indicate a larger difference between DDM1 and the real data. Similarly, in analyzing X-ray image generation, SW distance, MMD, and FID metrics suggest a high similarity between PGGAN-generated and real images, whereas C2ST indicates a strong difference. Thus, we want to highlight the importance of using multiple complementary distances for best results and understanding of model limitations.
@@ 5 Discussion
This work describes and characterizes four commonly applied sample-based distances representing different methodologies for defining statistical distance: Using low-dimensional projections (SW), obtaining a distance using classifiers (C2ST), using embeddings through kernels (MMD) or neural networks (FID). Despite their operational differences, they are all based on a fundamental concept: simplifying complex distributions into more manageable feature representations to facilitate comparison. Sliced distances effectively reduce multidimensional distributions to a set of one-dimensional distributions, where classical metrics are more easily applied or calculated. MMD uses kernels to (implicitly) project samples into a higher dimensional feature space, in which comparing mean values becomes more expressive. Classifier-based methods (C2ST) transform the task of distribution comparison into a classification problem; comparison is made by investigating how well a classifier can distinguish the distributions. Lastly, network-based distances, such as FID, explicitly map samples into a representative feature space and compare distributions directly within this space.
@@
Sliced distances stand out for their computational efficiency in evaluating distributional discrepancies. However, when distributions differ primarily in lower-dimensional subspaces, sliced distances might not detect these subtle differences without a large number of slices (see Fig. 8c). There are approaches to reduce this effect by considering other projections than simple linear slices, as described in Section 2.1. In our experiments, the metric did show convincing results and in contrast to the MMD, C2ST, and FID, SW distance does not require one to choose specific hyperparameters (for which results can differ drastically). Although currently not extensively used in literature for evaluation, this makes the SW distance efficient, scalable, and a consistent baseline for general distribution comparisons.Yet, this also makes it less flexible to adapt to specific features of interest. The Wasserstein and SW distances are not interpretable, and admit only biased sample-based estimates. This can be a limitation for some tasks. However, due to its computational efficiency and differentiability, the SW distance is commonly used as a loss function to train generative models, such as GANs (Deshpande et al., 2018, Wu et al., 2019), Autoencoders (Wu et al., 2019), nonparametric flows (Liutkus et al., 2019), normalizing flows (Dai & Seljak, 2021), and multi-layer perceptrons (Vetter et al., 2024). Although the majority of research on sliced distances focuses on sliced Wasserstein metrics, slicing other metrics is also possible. For a certain subset of choices, equivalence to MMDs can be established (Feydy et al., 2019, Kolouri et al., 2019, Hertrich et al., 2024).
@@
The Maximum Mean Discrepancy is a strong tool for comparing two groups of data by looking at their average values in a special feature space. The effectiveness of MMD largely depends on the kernel function chosen (implicitly representing the feature space), which affects how well it can spot differences between various types of data. Inappropriate kernel choice can leave the metric insensitive to subtle differences in the distribution (Gretton et al. 2012b, Sriperumbudur et al. 2009; see Fig. 8). The MMD can be estimated efficiently and is differentiable, and thus often used as a loss function for training generative models (Dziugaite et al., 2015, Arbel et al., 2019, Li et al., 2017, Bińkowski et al., 2018, Briol et al., 2019). Yet, a kernel must satisfy certain criteria, e.g., positive definiteness, making the design of new kernel functions challenging. Such constraints are relaxed for FID-like metrics, which focus on explicit representations of the embedding, whereas (kernel) MMD instead focuses on implicit representations. One advantage, however, is that the implicit embedding allows for infinite dimensional feature spaces (through characteristic kernel functions). These can be proven to be able to discriminate any two distinct distributions, something that is impossible through explicit representations used by the FID. Recently, Kübler et al. (2022a) proposed a method to estimate MMD via a witness function that determines MMD (Appendix A.2). This method is closely related to C2ST in that both estimate a discrepancy among distributions via a classifier (Kübler et al., 2022a, Section 5).
@@
Network-based approaches for evaluating distributions focus on the analysis of complex data, emphasizing the importance of capturing high-level, semantically meaningful features. These methods leverage neural networks to project data into a lower-dimensional, feature-rich space where traditional statistical distances can be applied more effectively. This is particularly important for tasks where the visual or semantic quality of the data is important, making them a popular choice for assessing generative models in domains such as image and text generation. The primary challenge lies in the design of suitable network architectures that can extract relevant features for accurate distribution comparison. Even more important than for the C2ST, this network must be well-established and shared which is a necessary but not sufficient criterion (Chong & Forsyth, 2020) to compare different results. While such well-established defaults exist for images (Szegedy et al., 2015, Radford et al., 2021), this is not the case for other domains. For example, the time series generation community did not yet establish a default, and embedding networks are either trained or chosen by the authors (Smith & Smith, 2020, Jeha et al., 2021). We demonstrated that the class-sensitivity of FID (Section 3.2) often leads to model collapse, as seen in GANs. However, it may not accurately reflect the overall image quality. For example, Betzalel et al. (2022), Kynkäänniemi et al. (2023), Jayasumana et al. (2023) found that relevant features sometimes can disagree with human judgment and that CLIP embeddings align more closely to what humans perceive as favorable or unfavorable. Yet, FID features have been shown to align much better with human perception than traditional metrics (Zhang et al., 2018). Recent developments in network-based approaches include the use of Central Kernel Alignment (CKA; Cortes et al. 2012) to compute the distance between network-embedded samples. CKA scores show considerable stability when evaluated with different choices of network architectures and layers (Yang et al., 2023). Another newly introduced metric, Mauve (Pillutla et al., 2021, Liu et al., 2021, Pillutla et al., 2023), can, for instance, be used to measure how close machine-generated text is to human language using an external language model to embed the samples from each distribution. This metric uses divergence frontiers to take into account the trade-off between quality and diversity when evaluating generative models.
@@
Ultimately, the choice of distance hinges on the nature of the data under consideration and the specific characteristics one aims to compare. Given a particular dataset and problem, it may be necessary to look beyond the distances discussed in this paper. For instance, in the realm of human-centric data such as images and audio, perceptual indistinguishability of distributions is crucial (Gerhard et al., 2013, Zhang et al., 2018). Time series data, with its inherent temporal structure, requires metrics that accommodate temporal shifts and variations without disproportionately penalizing minor discrepancies in timing, such as Dynamic Time Warping (Müller, 2007) or frequency-based methods (Hess et al., 2023). Additional recent works propose new distances based on fields such as topology (Barannikov et al. 2021).
In general, regardless of the specific use case, it is advisable to use multiple distance measures to obtain a comprehensive view, as relying on a single measure may lead to competing conclusions about the model under evaluation.
Throughout this paper, we have explained and analyzed four approaches to measuring statistical distance. While this represents only a small subset of all possible measures, we hope to have provided the foundational knowledge researchers need to find, understand, and interpret statistical distances relevant to their own scientific applications.
@@
Writing—Review & Editing: AD, AS, CS, FP, GM, JHM, JKL, JV, LH, LU, MG, MP, RG, SB, ST, ZS
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039; DeepHumanVision, FKZ: 031L0197B; Simalesam: FKZ 01|S21055A-B), the Carl Zeiss Foundation (Certification and Foundations of Safe Machine Learning Systems in Healthcare), the Else Kröner Fresenius Stiftung (Project ClinbrAIn), and the European Union (ERC, DeepCoMechTome, 101089288). ST was partially supported by Grant-in-Aid for Research Activity Start-up 23K19966, Grant-in-Aid for Early-Career Scientists 24K20750 and JST CREST JPMJCR2015. SB, MD, MG, JK, JL, GM, MP, RR, AS, ZS and JV are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
@@ A.1 Generative Models in Science
Table S1: Example generative models in science. This is not an exhaustive list regarding disciplines using generative models nor generative models used in the listed disciplines.
Discipline Model
Biology
[…]
@@ A.2 Details about Maximum Mean Discrepancy
Generally, $\mathcal{X}$ and $\mathcal{H}$ are defined as a topological space, and the reproducing kernel Hilbert space (RKHS), respectively, but readers can simply think of the Euclidean space $\mathbb{R}^{N}$ for the first examples in the main text.
@@
The definition of MMD can be rewritten through the notion kernel mean embedding. For given distribution $p(x)$, the kernel mean embedding $\mathbb{E}_{p(x)}[k(x,u)]\in\mathcal{H}$ is an element in RKHS that satisfies $\langle\mathbb{E}_{p(x)}[k(x,u)],f(u)\rangle_{\mathcal{H}}=\mathbb{E}_{p(x)}[f(x)]$ for any $f\in\mathcal{H}$ with argument $u\in\mathcal{X}$. The embedding $\mathbb{E}_{p(x)}[k(x,u)]$ is known to be determined uniquely if a corresponding kernel is bounded, i.e., $\|k(x,x^{\prime})\|_{\mathcal{H}}<\infty$ for any $x$. Then, as shown in Gretton et al. (2012a), $\mathsf{MMD}^{2}$ can be represented as
@@
A.3 Formal Definition of Wasserstein and Sliced-Wasserstein Distance
@@
The Sliced-Wasserstein distance is closely connected to the Radon transform (Helgason, 1980). We direct readers to Bonneel et al. (2015) for details. Here we provide a shorter definition following Definition 2.9 of Nadjahi (2021).
Suppose that $M\subset\mathbb{R}^{d}$, and denote by $\mathbb{S}^{d-1}=\{\theta\in\mathbb{R}^{d}\,:\,||\theta||_{2}=1\}$ be the unit sphere with respect to the Euclidean norm. Let $u^{*}:X\to\mathbb{R}$ be a the linear form given by $u^{*}(x)=\langle u,x\rangle$, and $q\in[1,+\infty)$. Then the Sliced-Wasserstein distance (of order $q$) is defined for any measures $\mu,\nu\in P_{q}(X)$ as
[…]
Where $\mathcal{U}$ is the uniform distribution on $\mathbb{S}^{d-1}$, and for any $u\in\mathbb{S}^{d-1}$, $u_{\#}^{*}$ denotes the push-forward operator of $u^{*}$.
As for the Wasserstein distance, the definition becomes more intuitive in the case where $\mu$ and $\nu$ admit the probability density functions ($p_{1}$ and $p_{2}$ respectively). In particular, the random projection $u$ are uniformly random vectors in $\mathbb{S}^{d-1}$. Therefore, projecting the distributions $p_{1}$ and $p_{2}$ onto $u$ induces one-dimensional distributions $p_{1}^{u}$ and $p_{2}^{u}$ with samples $u^{\mathsf{T}}x_{i}$, where $x_{1}\sim p_{1}$ and $x_{2}\sim p_{2}$. The Sliced-Wasserstein-$q$ distance can then be written as
$SW_{q}(p_{1},p_{2})=\left(\mathbb{E}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{q}^{q}(p_{1}^{u},p_{2}^{u})]\right)^{1/q}.$ (7)
A.4 Dependence of SW Distance on Number of Projections
Figure S1: The SW distance estimate is not strongly sensitive to the number of random projections. We compare the SW distance estimate for the $\{10,100,1000\}$-dimensional Gaussian task with 1 shifted dimension (Sec. 3) as we increase the number of random projections used in the estimation. As the number of projections increases, the variance of the SW distance estimate decreases, but across all dimensionalities considered, the SW distance estimate has converged by 100 random projections.
All SW distance experiments in the main text were performed with the SW distance approximate with 100 random projections to approximate the expectation Eq. (7). Here, we additionally show the dependence of the SW distance approximation with finite projections on the $d$-dim Gaussian example (see Sec.±3), for $d\in\{10,100,1000\}$ (Fig S1). While the sample-based approximation to the analytic 1-dimensional Wasserstein distance is biased (Sec. 2.1), the Monte Carlo approximation to the expectation Eq. (7) is an unbiased estimate of the sample-based Wasserstein distance. We observe that for very high (1000)-dimensional distributions, the SW distance estimate converges quickly, and the choice of 100 random projections for the computation of the SW distance is appropriate.
A.5 C2ST scores below 0.5